# Practical Poisoning Attacks against Retrieval-Augmented Generation

### v1

@@ 4.4 CorruptRAG-AK
To create a generalized method for crafting adversarial knowledge across targeted queries, we use an LLM (e.g., GPT4o-mini) to generate this adversarial knowledge. Notably, the LLM used for crafting may differ from the one in the RAG system. In CorruptRAG-AK, the attacker first generates $p_{i}^{h}$ as in CorruptRAG-AS, then uses few-shot learning to craft a prompt guiding the LLM to refine $p_{i}^{h}$. The prompt is shown in Appendix A, where the parameter $V$ specifies the length of $p_{i}^{h}$. Since a securely aligned LLM may refine $q_{i}$ in favor of the correct answer $C_{i}$ instead of the targeted answer $A_{i}$, once $p_{i}^{h}$ is refined, we use it as context for the LLM to generate an answer for $q_{i}$. If the response does not match $A_{i}$, we re-prompt the LLM to refine $p_{i}^{h}$ until success or until reaching the maximum number of attempts, $L$. We show serval examples of CorruptRAG-AS and CorruptRAG-AK in Tables 21-22 in Appendix.
Table 1: Results of various attacks across three datasets.
Attack Metrics GPT-3.5-turbo GPT-4o-mini GPT-4o GPT-4-turbo
PoisonedRAG (Black-Box) ASR 0.54 0.69 0.52 0.58
Recall 0.99
F1-score 0.33
PoisonedRAG (White-Box) ASR 0.75 0.67 0.56 0.57
Recall 1.00
F1-score 0.33
PIA ASR 0.76 0.85 0.67 0.78
Recall 0.90
F1-score 0.30
CPA ASR 0.06 0.02 0.02 0.03
Recall 1.00
F1-score 0.33
CorruptRAG-AS ASR 0.90 0.97 0.89 0.94
Recall 0.98
F1-score 0.33
CorruptRAG-AK ASR 0.94 0.95 0.85 0.93
Recall 0.98
F1-score 0.33
Attack Metrics GPT-3.5-turbo GPT-4o-mini GPT-4o GPT-4-turbo
PoisonedRAG (Black-Box) ASR 0.60 0.83 0.67 0.77
Recall 1.00
F1-score 0.33
PoisonedRAG (White-Box) ASR 0.57 0.66 0.70 0.71
Recall 1.00
F1-score 0.33
PIA ASR 0.88 0.95 0.78 0.93
Recall 1.00
F1-score 0.33
CPA ASR 0.04 0.01 0.01 0.01
Recall 1.00
F1-score 0.33
CorruptRAG-AS ASR 0.92 0.98 0.84 0.97
Recall 1.00
F1-score 0.33
CorruptRAG-AK ASR 0.94 0.97 0.89 0.94
Recall 1.00
F1-score 0.33
Attack Metrics GPT-3.5-turbo GPT-4o-mini GPT-4o GPT-4-turbo
PoisonedRAG (Black-Box) ASR 0.55 0.69 0.61 0.57
Recall 0.97
F1-score 0.32
PoisonedRAG (White-Box) ASR 0.48 0.59 0.55 0.54
Recall 0.98
F1-score 0.33
PIA ASR 0.72 0.87 0.64 0.77
Recall 0.89
F1-score 0.30
CPA ASR 0.06 0.10 0.07 0.07
Recall 0.99
F1-score 0.33
CorruptRAG-AS ASR 0.87 0.92 0.85 0.94
Recall 0.95
F1-score 0.32
CorruptRAG-AK ASR 0.86 0.96 0.88 0.92
Recall 0.99
F1-score 0.33
Table 2: Results of different retrievers on NQ dataset.
Attacks Metrics Contriever Contriever-ms ANCE
@@
Table 3: Results of different similarity metrics on NQ dataset.
Attack Metrics Dot Product Cosine Similarity
@@
Table 4: Results of concatenation order of $p^{s}_{i}$ and $p^{h}_{i}$ on NQ dataset.
Attack Metrics $p^{s}_{i}\oplus p^{h}_{i}$ $p^{h}_{i}\oplus p^{s}_{i}$
CorruptRAG-AS ASR 0.97 0.78
Recall 0.98 0.98

### v2

@@ 4.4. CorruptRAG-AK
To create a generalized method for crafting adversarial knowledge across targeted queries, we use an LLM (e.g., GPT4o-mini) to generate this adversarial knowledge. Notably, the LLM used for crafting may differ from the one in the RAG system. In CorruptRAG-AK, the attacker first generates $p_{i}^{h}$ as in CorruptRAG-AS, then uses few-shot learning to craft a prompt guiding the LLM to refine $p_{i}^{h}$. The prompt is shown as follow, where the parameter $V$ specifies the length of $p_{i}^{h}$.
Since a securely aligned LLM may refine $q_{i}$ in favor of the correct answer $C_{i}$ instead of the targeted answer $A_{i}$, once $p_{i}^{h}$ is refined, we use it as context for the LLM to generate an answer for $q_{i}$. If the response does not match $A_{i}$, we re-prompt the LLM to refine $p_{i}^{h}$ until success or until reaching the maximum number of attempts, $L$. Tables 14 and 15 in Appendix show examples of poisoned texts crafted by CorruptRAG (CorruptRAG-AS and CorruptRAG-AK) on NQ and MS-MARCO datasets.
5. Experiments
5.1. Experimental Setup
Table 1. Statistics of three datasets.
Datasets #Texts #Queries
NQ 2,681,468 3,452
HotpotQA 5,233,329 7,405
MS-MARCO 8,841,823 6,980
5.1.1. Datasets
In our experiments, we utilize three large-scale datasets from the Beir benchmark (Thakur et al., 2021) related to the information retrieval task of English: Natural Questions (NQ) (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and MS-MARCO (Nguyen et al., 2016). The statistics of the three datasets are summarized in Table 1.
Natural Questions (NQ) (Kwiatkowski et al., 2019): The NQ dataset is derived from Wikipedia and includes 2,681,468 texts. Its test set consists of 3,452 queries which are sampled from Google search history.
HotpotQA (Yang et al., 2018):  The knowledge database for HotpotQA is also collected from Wikipedia and contains 5,233,329 texts. This database comprises 7,405 queries.
MS-MARCO (Nguyen et al., 2016): The knowledge database of MS-MARCO is sourced from web documents retrieved by Bing and comprises 8,841,823 texts and 6,980 queries.
Note that in the absence of attacks, the RAG system produces highly accurate responses to targeted queries, with accuracy rates of 81% on NQ, 80% on HotpotQA, and 84% on MS-MARCO. These results are consistent with those reported in PoisonedRAG (Zou et al., 2025).
Table 2. Attack results on three datasets. Higher ($\uparrow$) ASR, Recall, and F1-score indicate better attack performance. Note that because only a single poisoned text is injected for each targeted query, the maximum attainable F1-score in our default setting (e.g., when the top 5 texts are retrieved) is 0.33.
Datasets Attacks Metrics GPT-3.5-turbo GPT-4o-mini GPT-4o GPT-4-turbo
NQ PoisonedRAG (Black-box) ASR ($\uparrow$) 0.54 0.69 0.52 0.58
Recall ($\uparrow$) 0.99
F1-score ($\uparrow$) 0.33
PoisonedRAG (White-box) ASR ($\uparrow$) 0.75 0.67 0.56 0.57
Recall ($\uparrow$) 1.00
F1-score ($\uparrow$) 0.33
PIA ASR ($\uparrow$) 0.76 0.85 0.67 0.78
Recall ($\uparrow$) 0.90
F1-score ($\uparrow$) 0.30
CPA ASR ($\uparrow$) 0.06 0.02 0.02 0.03
Recall ($\uparrow$) 1.00
F1-score ($\uparrow$) 0.33
CorruptRAG-AS ASR ($\uparrow$) 0.90 0.97 0.89 0.94
Recall ($\uparrow$) 0.98
F1-score ($\uparrow$) 0.33
CorruptRAG-AK ASR ($\uparrow$) 0.94 0.95 0.85 0.93
Recall ($\uparrow$) 0.98
F1-score ($\uparrow$) 0.33
HotpotQA PoisonedRAG (Black-box) ASR ($\uparrow$) 0.60 0.83 0.67 0.77
Recall ($\uparrow$) 1.00
F1-score ($\uparrow$) 0.33
PoisonedRAG (White-box) ASR ($\uparrow$) 0.57 0.66 0.70 0.71
Recall ($\uparrow$) 1.00
F1-score ($\uparrow$) 0.33
PIA ASR ($\uparrow$) 0.88 0.95 0.78 0.93
Recall ($\uparrow$) 1.00
F1-score ($\uparrow$) 0.33
CPA ASR ($\uparrow$) 0.04 0.01 0.01 0.01
Recall ($\uparrow$) 1.00
F1-score ($\uparrow$) 0.33
CorruptRAG-AS ASR ($\uparrow$) 0.92 0.98 0.84 0.97
Recall ($\uparrow$) 1.00
F1-score ($\uparrow$) 0.33
CorruptRAG-AK ASR ($\uparrow$) 0.94 0.97 0.89 0.94
Recall ($\uparrow$) 1.00
F1-score ($\uparrow$) 0.33
MS-MARCO PoisonedRAG (Black-box) ASR ($\uparrow$) 0.55 0.69 0.61 0.57
Recall ($\uparrow$) 0.97
F1-score ($\uparrow$) 0.32
PoisonedRAG (White-box) ASR ($\uparrow$) 0.48 0.59 0.55 0.54
Recall ($\uparrow$) 0.98
F1-score ($\uparrow$) 0.33
PIA ASR ($\uparrow$) 0.72 0.87 0.64 0.77
Recall ($\uparrow$) 0.89
F1-score ($\uparrow$) 0.30
CPA ASR ($\uparrow$) 0.06 0.10 0.07 0.07
Recall ($\uparrow$) 0.99
F1-score ($\uparrow$) 0.33
CorruptRAG-AS ASR ($\uparrow$) 0.87 0.92 0.85 0.94
Recall ($\uparrow$) 0.95
F1-score ($\uparrow$) 0.32
CorruptRAG-AK ASR ($\uparrow$) 0.86 0.96 0.88 0.92
Recall ($\uparrow$) 0.99
F1-score ($\uparrow$) 0.33
5.1.2. Comparison of Attacks
  • Black-box setting: The attacker has no access to the parameters of the LLM and the retriever, and only uses an LLM to craft the poisoned text for the targeted queries.
  • White-box setting:  The attacker knows the retriever’s parameters, enabling the further optimization of the poisoned text to maximize its similarity with the targeted query.
Prompt injection attack (PIA) (Perez and Ribeiro, 2022; Zou et al., 2025): This attack was initially designed for LLMs and later adapted to RAG systems by (Zou et al., 2025). The attacker crafts the poisoned texts by concatenating the targeted query with a malicious prompt that instruct the LLM to generate the targeted answer.
5.1.3. Evaluation Metrics
We consider three metrics: attack success rate (ASR), Recall, and F1-score.
Attack success rate (ASR): ASR is defined as the proportion of queries that yield RAG outputs matching the targeted answers among all targeted queries. We employ an accurate LLM judgment method, which leverages GPT-4o-mini to evaluate the consistency between the RAG output and the targeted answer.
Recall: Recall is defined as the proportion of successfully retrieved poisoned texts within the top-$N$ among all injected poisoned texts for each targeted query. Since we only inject one poisoned text per targeted query across all attacks, Recall can be calculated as the proportion of targeted queries where the poisoned text appears within the top-$N$.
F1-score: We first introduce the Precision, which is the proportion of poisoned texts among the retrieved top-$N$ texts for the targeted query. Then, F1-score is defined as $\text{F1-score}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$.
Higher ASR, Recall, and F1-score signify stronger attack performance. Note that since only one poisoned text is injected per targeted query, the maximum achievable F1-score is constrained to $\frac{2}{N+1}$. For instance, with $N=5$, the highest possible F1-score is 0.33.
5.1.4. Parameter Setting
Following (Zou et al., 2025), we randomly select 100 closed-ended queries per dataset as targeted queries and employ an LLM, such as GPT-4o-mini, to generate random answers different from the correct ones as targeted answers. For the RAG system, we set $N=5$ (i.e., top-5 relevant texts are retrieved by the retriever), using GPT-4o-mini as the LLM, Contriever (Izacard et al., 2021) as the retriever and the dot product as the similarity metric. For CorruptRAG-AK, we use GPT-4o-mini to craft $p^{h}_{i}$ with $V=30$ and $L=5$. Note that, for a fair comparison, across all attacks (PoisonedRAG (Zou et al., 2025), PIA (Perez and Ribeiro, 2022; Zou et al., 2025), CPA (Zhong et al., 2023), and our proposed CorruptRAG), we inject one poisoned text per targeted query. All experiments were conducted on a server equipped with an Intel Gold 6248R CPU and four NVIDIA 3090 GPUs. Each experiment was repeated 10 times, and the average results were reported.
Table 3. Results of PoisonedRAG on the NQ dataset when the attacker injects five poisoned texts per targeted query.
Attacks ASR Recall F1-score
PoisonedRAG (Black-box) 0.89 0.96 0.96
PoisonedRAG (White-box) 0.95 1.00 1.00
Table 4. Price (USD) of crafting poisoned texts for each query.
Datasets CorruptRAG-AS CorruptRAG-AK
NQ 0.0000 0.0001
HotpotQA 0.0000 0.0001
MS-MARCO 0.0000 0.0001
5.2. Experimental Results
5.2.1. Main Results
Our CorruptRAG attacks outperform all baseline attacks:  We conduct a comprehensive evaluation of our CorruptRAG attacks and several baseline attacks across three diverse datasets. This evaluation utilizes RAG systems powered by a range of prominent LLMs, including GPT-3.5-turbo, GPT-4o-mini, GPT-4o, and GPT-4-turbo. The results, presented in Table 2, demonstrate the superior performance of our CorruptRAG attacks, decisively outperforming all baseline attacks, particularly evidenced by the significantly higher ASR.
A key observation highlights a substantial degradation in the effectiveness of baseline attacks when limited to injecting only a single poisoned text per targeted query. This performance drop stems from the fact that their approaches to crafting poisoned texts do not account for the stringent constraint on the number of poisoned texts per targeted query (as detailed in Section 4.1). Consequently, the manipulative influence exerted by a single poisoned text crafted by these baseline attacks is often limited, allowing the LLM to predominantly rely on the correctly relevant texts and ultimately produce the correct answer in most instances. In stark contrast, our CorruptRAG attacks are fundamentally designed to maximize attack potency under such constraints. We explicitly optimizes each poisoned text to ensure high individual effectiveness. This inherent focus on maximizing the impact of a single instance explains why our CorruptRAG attacks consistently achieve high ASRs.
Practical limitations of PoisonedRAG: As shown in Table 2, PoisonedRAG performs poorly when an attacker can insert only one poisoned text per targeted query. To probe the limitation of PoisonedRAG, we also evaluate a setting where the attacker in PoisonedRAG injects multiple poisoned texts per query (e.g., five). Results on the NQ dataset reported in Table 3 match those in the original PoisonedRAG paper. From Section 4.1, most queries in NQ have at most three relevant texts in the knowledge database. Consequently, inserting five poisoned texts per query makes the poisoned content exceed the number of relevant texts. Although this amplifies the attack effect, it represents an unrealistic operating point: such dense poisoning would overwhelm the knowledge database and is likely to trigger integrity or anomaly detection, making PoisonedRAG impractical in real deployments.
Our CorruptRAG attacks are cost-effective: We analyze the monetary cost of our CorruptRAG attacks by measuring the LLM API expenses for crafting poisoned text per targeted query. Based on the official pricing for GPT-4o-mini (\$0.15 USD per 1 million input tokens and \$0.60 USD per 1 million output tokens (36)), Table 4 presents the average per-query cost incurred by our attacks across the three datasets. We highlight two key observations. Firstly, CorruptRAG-AS incurs absolutely no cost, as it generates highly effective poisoned text directly by populating predefined templates with the correct answer and intended targeted answer of the targeted query, completely bypassing the need for LLM API calls during generation. Secondly, the API cost associated with CorruptRAG-AK, while non-zero, is remarkably low, averaging approximately \$0.0001 USD per query, making it practically negligible. These results powerfully demonstrate the value of optimizing for single-text effectiveness under the constraint of limited poisoned texts per query (detailed in Section 4.1), a core principle of our attack design that drastically reduces costs. This characteristic of being exceptionally cheap (even zero-cost) to implement renders our attacks highly practical.
Table 5. Results of different retrievers.
Datasets Attacks Metrics Contriever Contriever-ms ANCE
NQ PoisonedRAG (Black-box) ASR 0.69 0.55 0.47
Recall 0.99 1.00 0.99
F1-score 0.33 0.33 0.33
PoisonedRAG (White-box) ASR 0.67 0.53 0.50
Recall 1.00 1.00 1.00
F1-score 0.33 0.33 0.33
PIA ASR 0.85 0.85 0.87
Recall 0.90 0.98 1.00
F1-score 0.30 0.33 0.33
CPA ASR 0.02 0.02 0.02
Recall 1.00 1.00 0.97
F1-score 0.33 0.33 0.32
@@ 5.2.1. Main Results
MS-MARCO PoisonedRAG (Black-box) ASR 0.69 0.47 0.44
Recall 0.97 0.99 1.00
F1-score 0.32 0.33 0.33
PoisonedRAG (White-box) ASR 0.59 0.32 0.37
Recall 0.98 1.00 0.99
F1-score 0.33 0.33 0.33
PIA ASR 0.87 0.83 0.83
Recall 0.89 0.98 1.00
F1-score 0.30 0.33 0.33
CPA ASR 0.10 0.09 0.09
Recall 0.99 1.00 0.94
F1-score 0.33 0.33 0.31
CorruptRAG-AS ASR 0.92 0.83 0.87
Recall 0.95 0.99 0.99
F1-score 0.32 0.33 0.33
CorruptRAG-AK ASR 0.96 0.87 0.84
Recall 0.99 0.99 0.98
F1-score 0.33 0.33 0.33
Table 6. Results of different similarity metrics.
Datasets Attacks Metrics Dot Product Cosine Similarity
NQ PoisonedRAG (Black-box) ASR 0.69 0.8
Recall 0.99 1.00
F1-score 0.33 0.33
PoisonedRAG (White-box) ASR 0.67 0.76
Recall 1.00 0.98
F1-score 0.33 0.33
PIA ASR 0.85 0.84
Recall 0.90 0.92
F1-score 0.30 0.31
CPA ASR 0.02 0.01
Recall 1.00 1.00
F1-score 0.33 0.33
@@
MS-MARCO PoisonedRAG (Black-box) ASR 0.69 0.68
Recall 0.97 0.99
F1-score 0.32 0.33
PoisonedRAG (White-box) ASR 0.59 0.62
Recall 0.98 0.88
F1-score 0.33 0.29
PIA ASR 0.87 0.80
Recall 0.89 0.86
F1-score 0.30 0.29
CPA ASR 0.10 0.04
Recall 0.99 1.00